OpenStudy (anonymous):

Calculate the equation of the line between the points (-6,-1) and (2,-1)

OpenStudy (anonymous):

use this plug in values and simplify

OpenStudy (anonymous):

\[\frac{ (y-y1) }{ (x-x1) }=\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome/..